OpenStudy (anonymous):

howdy can some help me with, the midpoint of uv is (5 -11) the coordinates of one end point is (3,5) find the coordinates of end point v

OpenStudy (anonymous):



OpenStudy (anonymous):

Have you got it ?

OpenStudy (anonymous):

No i still dont understand can you help solve it in steps?

OpenStudy (anonymous):

\[5=\frac{ 3+x }{2 },3+x=10,x=10-3=7\] \[-11=\frac{ 5+y }{2 },5+y=-22,y=?\]

OpenStudy (anonymous):

would it be 22-5?

OpenStudy (anonymous):

y=-22-5=-27

OpenStudy (anonymous):

end point is (7,-27)

OpenStudy (anonymous):

Thank you kind sir you really are a kind soul

OpenStudy (anonymous):

yw